OpenStudy (anonymous):

What are the x-intercept(s) of the graph of y + 6 = x2 − x? (−2, 0) and (3, 0) (−2, 0) and (−3, 0) (2, 0) and (−3, 0) (2, 0) and (3, 0)

OpenStudy (saifoo.khan):

When x-intercepts, y = 0. y + 6 = x2 − x 0 + 6 = x2 - x x^2 - x - 6 = 0 Now here's a quadratic equation. Can you solve from here?

OpenStudy (anonymous):

so it's (2,0) and (-3,0)?